OpenStudy (jiteshmeghwal9):

A company reduced the price of an article by 5% . By how much a retailer must increase the new price so as to sell the article at the old price ?

OpenStudy (shubhamsrg):

let original be x after decrease, new price = 0.95 x if we take it back to x, increase = (x - 0.95x)/0.95 x * 100 = 0.05/0.95 * 100 just simplify..

OpenStudy (jiteshmeghwal9):

 Thanx a lot bro :)